Assuming that the earth re-emits all the radiations it receives from the sum, determine
its temperature, (For earth, E = 1.39 × 103
w/m2 , Stefan’s constant. = σ = 5.735 × 10–8
w/m2
k
4
)
1
Expert's answer
2019-10-24T09:36:20-0400

Determining the temperature

We need to find the temperature


Solution:


We have formula,

"T = [\frac {E \times (1 - Albedo)} {4\sigma}]^\frac {1}{4}"

Here,

Albedo = 0.3


Solar constant =

"E = 1.39 \times 10^3 \frac {w} {m^2}"



Stefan’s constant. =

"\sigma =5.735 × 10^{–8}\frac {w} {m^2}"

Plug the values in the formula



"T = [\frac {1.39 \times 10^3 \times (1 - 0.3)} {4\times 5.735 × 10^{–8} }]^\frac {1}{4}"

"T = 255.2 \space k"


Answer:

"T = 255.2 \space k"